OpenStudy (anonymous):

Find the four arithmetic means between 12 and -13. 6, -1, -3, -8 6, 2, -3, -8 7, -1, -3, -8 7, 2, -3, -8

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (akashdeepdeb):

To solve these kinds of questions you need to know some stuff. You will have 12 _ _ _ _ -13 a = 12 An = -13 n = 6 So we have to find d [Common difference] :D An = a + (n-1)d -13 = 12 + (5)d d = -5 Now I hope you know how to go about this... :D Understood this? :)

OpenStudy (anonymous):

thw ques is if a,b,c in A.P. and x is the arthmetic mean of a and b ,y is the artmetic mean of b and c then how to prove that a,x,b,y,c is in A.P

OpenStudy (akashdeepdeb):

thw?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

srry it's the

OpenStudy (akashdeepdeb):

Did you get the solution to the previous one??

OpenStudy (anonymous):

no

OpenStudy (akashdeepdeb):

What did you not get in that?

OpenStudy (akashdeepdeb):

Don't jump to the next question.. Finish one and then do.

OpenStudy (anonymous):

the ques is if a,b,c in A.P. and x is the arthmetic mean of a and b ,y is the artmetic mean of b and c then how to prove that a,x,b,y,c is in A.P

OpenStudy (akashdeepdeb):

WHAT? ARE YOU EVEN LISTENING/LOOKING AT WHAT I TYPED??????

OpenStudy (anonymous):

this is my first day and i don't know if have done something wrong but still i have not got my ans

OpenStudy (akashdeepdeb):

No problemo! :) Have you done AP bfore? If you haven't look up some websites/READ, If you have ask what you did not understand. :)

OpenStudy (anonymous):

ya i have learned ap but where is ur ans i can't see here

OpenStudy (akashdeepdeb):

We cannot give answers here... We can help you but it is your question and you have to find it out!! :D I can guide you a 100 times if you want!! :D

OpenStudy (anonymous):

ok then just give a hint about the solution of this question

OpenStudy (akashdeepdeb):

You will have 12 _ _ _ _ -13 a = 12 An = -13 n = 6 So we have to find d [Common difference] :D An = a + (n-1)d -13 = 12 + (5)d d = -5 COMMON DIFFERENCE = -5 Second term = a + d = 12 - 5 = 7 Third term = a+2d = 12-10 = 2 Understood this? :)

OpenStudy (anonymous):

YA I KNW VERY WELL about this

OpenStudy (anonymous):

but my question is different

OpenStudy (akashdeepdeb):

You have posted 2 different questions in the same question. Did you understand and get the answer to the first question?